Consider u(x_{1},x_{2})=x_{1}^{0.5}x_{2}^{0.5} ,I=600,p_{1}=25,p_{2}=1. If p_{1} decreases from $25 to $9, what is the compensating variation? Enter a number only, round to two decimals. If money needs to be taken away from the consumer include a negative sign.8.2 Homework • Unanswered Consider u(x1, x2) = x9.5x9.5,1 = 600,p1 = 25,P2 = 1. If P1 decreases from $25 to $9, what is the c
Answer #1

0.5 0.5 u - mu, = n, du ona na -0.5. n.os na My2 = 0u - 0.5nos neos. ang At optimun mu - P - a mur 0.5m, os nos – 0.5 nos noU = I 2 RP2) os I = 2(8, P2J. U Now, P, = 9, P2=1, u= 20.7846 I= 2(9x) 20.7846 = 124.7076 But given income is 600 So compensa